OpenStudy (anonymous):

Please Help! The volume of a sample of hydrogen is 798 mL and it exerts 621 mm Hg pressure at 5.00ºC. What volume does it occupy at standard temperature and pressure?

OpenStudy (somy):

\[\frac{ P_1\times V_1 }{ T_1 } = \frac{ P_2 \times V_2 }{ T_2 }\] P1 = 621 mm Hg - convert to atm V1 = 798 mL - convert to L T1 = 5.00ºC convert to Kelvin by adding 273 to the Celsius value now they gave condition- standard temperature and pressure so: P2 = 1 atm T2 = 273 Kelvin V2 = Unknown so now you can put the values into this formula and get final answer hope it helps :))